# Delin Qu∗1, Haoming Song∗1, Qizhi Chen∗1, Yuanqi Yao1, Xinyi Ye1, Yan Ding1, Zhigang Wang1 Jiayuan Gu3, Bin Zhao1, Dong Wang1, Xuelong Li12 1Shanghai AI Laboratory, 2TeleAI, 3ShanghaiTech

### v4

@@ IV-B Adapting to New Robot Setups
Evaluation Setup and Comparisons. We present the evaluation of SpatialVLA on the LIBERO simulation benchmark [35], which consists of a set of diverse robotic manipulation tasks in simulated environments. Following OpenVLA [29], we conduct experiments on four task suites, each comprising 10 tasks with 50 human-teleoperated demonstrations. These suites evaluate the model’s understanding of spatial relationships (LIBERO-Spatial), object types (LIBERO-Object), task-oriented behaviors (LIBERO-Goal), and its ability to generalize to long-horizon tasks with diverse objects, layouts, and goals (LIBERO-Long). We compare our approach to several generalist manipulation policy methods, including Diffusion Policy [12], Octo [46], OpenVLA [29], and TraceVLA [65]. SpatialVLA is fine-tuned on the corresponding dataset for 200 epochs using LoRA ($r=32$, $\alpha=32$), which incorporates spatial embedding adaption in Section. III-B from new Gaussian distribution.
@@
To facilitate a more comprehensive evaluation, 13 Franka tasks are established to validate the model’s manipulation performance, as shown in Figure. 5. The evaluation consists of three setups: Single Task, which includes four basic tasks: pick, place, push, and close; Instruction Following, which involves manipulating different objects in the same scene based on language instructions; and Multi-tasks, which involves training on a mixture of all four single-task data and tested on these tasks. We compare SpatialVLA with mainstream policies, including Diffusion Policy, Octo, and OpenVLA. More details can be found in the Section. -D.
@@
Evaluation Results. Table. III present the LIBERO [35] experimental results. Notably, we observe that SpatialVLA can be effectively adapted to tasks in the LIBERO environments, as it obtains the highest average success rate of 78.1% and the first rank across all the policies. In particular, SpatialVLA achieves a remarkable 88.2% success rate on the LIBERO-Spatial task, which consists of different object layouts, demonstrating the model’s strong understanding of spatial relationships. In most tasks, SpatialVLA outperforms the state-of-the-art generalist manipulation policies but struggles with long-horizon tasks in LIBERO-Long, due to the lack of architecture design for long-horizon observation.
@@
Figure. 5 summarizes the results of the Franka robot fine-tuning evaluation. In single-task tests, SpatialVLA and Diffusion Policy show similar accuracy (82% vs 81%), outperforming OpenVLA and Octo. However, in the instruction following tasks, SpatialVLA improves by +12% over OpenVLA, while Diffusion Policy struggles with a 26% success rate. In multi-tasks, SpatialVLA leverages its pre-training on OXE and 3D perception capabilities to achieve a 57% accuracy rate, surpassing other generalist policies. In summary, SpatialVLA demonstrates its versatility as a generalist robot control policy, achieving better performance across various tasks, and can be effectively used as an initialization for new robot tuning.
@@ IV-C Evaluating Spatial Understanding Capability
Evaluation Setup and Comparisons. As shown in Figures. 6 and III, we evaluate the spatial understanding capabilities of SpatialVLA through on three robot setups: Franka Robot fine-tuning, WidowX Robot zero-shot, and Libero-Spatial fine-tuning. The tasks exhibit varying spatial complexities, with the Franka task involving prompt understanding (e.g., #1 place plush toy closest to robot on car), the WidowX task featuring explicit height changes (e.g., #2 put green cup on the pink cloth), and the LIBERO-Spatial task involving object layout variations. Seven mainstream policies, namely Diffusion Policy, Octo, RT-1-X, OpenVLA, TraceVLA, and RoboVLM, are employed for comparison.
Evaluation Results. Compared to existing policies, SpatialVLA shows superior spatial understanding, achieving 73% accuracy in Franka task #1, which involves spatial prompts, and significantly improving manipulation capabilities for complex positional changes in the out-of-distribution WidowX Zero-shot tasks #2-4. Similar results are observed in the LIBERO-Spatial task suite (88.2% success rate). Policies like Octo, Diffusion Policy, and OpenVLA, which lack integrated depth information, face significant challenges in adapting to spatial layout changes, yielding a success rate consistently lower than 50%. Consequently, we suggest integrating 3D information (Section. III), including depth or point cloud, into the VLA framework to improve the model’s adaptability and robustness in spatial layout variations.
@@
$[1]$. SpatialVLA 81.6% 70.7% 79.2% 85.4% 41.7% 33.3% 91.7% 87.5%
$[2]$. $\sim$ linear 256 bins 40.7% 19.0% 47.1% 52.9% 41.7% 33.3% 87.5% 70.8%
$[3]$. $\sim$ uniform distribution 77.9% 28.0% 64.2% 55.0% 45.8% 12.5% 79.2% 54.2%
$[4]$. $\sim$ resolution 1026 74.4% 67.3% 59.1% 54.2% 45.8% 25.0% 66.7% 54.2%
$[5]$. -- ego3d encoding 68.9% 70.3% 66.7% 62.0% 54.2% 12.5% 75.0% 37.5%
@@ IV-D Ablations on Design Decisions
Pre-training in Mixture Dataset. The pre-training ablations in Table. IV are conducted on a mixture dataset that combines Google Fractal [6] and BridgeData V2 [59]. All the models are trained from scratch on 8 A100 GPUs with 128 batch size for 120k steps. We select four tasks from the SimplerEnv benchmark [34], namely ”Pick Coke Can” and ”Move Near” on the Google Robot, as well as ”Put Carrot on Plate” and ”Put Eggplant in Yellow Basket” on the WidowX Robot, to dissect the model’s component-wise performance.
@@
Compared to 1026-resolution action grids (#1 v.s. #4), where $\text{M}_{\text{trans}}=\text{M}_{\text{trans}}=512,\text{M}_{\text{grip}}=2$, SpatialVLA with 8194-resolution action grids ($\text{M}_{\text{trans}}=\text{M}_{\text{trans}}=4096,\text{M}_{\text{grip}}=2$) achieves significant performance boosts, particularly in ”move near” and ”put eggplant in yellow basket” tasks, with success rate increments of +31.2% and +33.3%. Additionally, we find that lower-resolution models tend to learn smaller actions, causing slow motion issues, and high-resolution models exhibit improved transfer performance in the fine-tuning stage.
According to the ablation results (#1 v.s. #5), the proposed egocentric 3D position encoding (ego3d), incorporating 3D point cloud features, helps the model overcome varied lighting, color, textures, and camera poses, yielding stronger generalizability in diverse manipulation scenarios. Models w/o ego3d suffer a significant performance drop in variant aggregation, from 81.6% and 79.2% to 68.9% and 66.7%, due to their inability to adapt to scene changes. During pre-training, we also observe from (#1 v.s. #6) that freezing the language embedding and sharing a trainable spatial embedding helps to improve the model’s manipulation capabilities, which is also beneficial for faster training and instruction following.
Post-training in Domain Dataset. We conduct post-training ablations in Table. V, separately fine-tuning on large-scale datasets Google Fractal and BridgeData V2 and BridgeData V2, and comparing full fine-tuning and LoRA-tuning on the small-scale LIBERO datasets [35]. The spatial embedding adaption denotes partitioning spatial grids from the new dataset Gaussian distribution and updating the spatial feature embedding with the grids.
On large-scale datasets (#1 v.s. #2), models fine-tuned with spatial embedding adaptation yield marginal gains of +2.9% in visual matching on Move Near), as the large-scale), as the large-scale dataset distribution closely matches the pre-training distribution, allowing the model to learn fine-grained features thereby limiting the benefits of the adaption. While, on the LIBERO small dataset tasks (#4 v.s. #5), initializing the feature grid with the new distribution boosts model performance by +4.6%, +5.1%, +2.2%, and +5.4% on LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long, respectively. As shown in Figure. 7, feature adaptation from the new distribution aligns pre-trained spatial features with the target fine-tuned model, improving initialization and accelerating convergence. Moreover, LoRA fine-tuning outperforms full-parameter fine-tuning on small dataset tasks (#3 v.s. #4), making LoRA the preferred method for small datasets.

### v5

@@ IV-B Adapting to New Robot Setups
Evaluation Setup and Comparisons. We present the evaluation of SpatialVLA on the LIBERO simulation benchmark [36], which consists of a set of diverse robotic manipulation tasks in simulated environments. Following OpenVLA [30], we conduct experiments on four task suites, each comprising 10 tasks with 50 human-teleoperated demonstrations. These suites evaluate the model’s understanding of spatial relationships (LIBERO-Spatial), object types (LIBERO-Object), task-oriented behaviors (LIBERO-Goal), and its ability to generalize to long-horizon tasks with diverse objects, layouts, and goals (LIBERO-Long). We compare our approach to several generalist manipulation policy methods, including Diffusion Policy [12], Octo [48], OpenVLA [30], and TraceVLA [71]. SpatialVLA is fine-tuned on the corresponding dataset for 200 epochs using LoRA ($r=32$, $\alpha=32$), which incorporates spatial embedding adaption in Sec. III-B from new Gaussian distribution.
@@
To facilitate a more comprehensive evaluation, 13 Franka tasks are established to validate the model’s manipulation performance, as shown in Fig. 5. The evaluation consists of three setups: Single Task, which includes four basic tasks: pick, place, push, and close; Instruction Following, which involves manipulating different objects in the same scene based on language instructions; and Multi-tasks, which involves training on a mixture of all four single-task data and tested on these tasks. We compare SpatialVLA with mainstream policies, including Diffusion Policy, Octo, and OpenVLA. More details can be found in the Sec. -E.
@@
Evaluation Results. Tab. III present the LIBERO [36] experimental results. Notably, we observe that SpatialVLA can be effectively adapted to tasks in the LIBERO environments, as it obtains the highest average success rate of 78.1% and the first rank across all the policies. In particular, SpatialVLA achieves a remarkable 88.2% success rate on the LIBERO-Spatial task, which consists of different object layouts, demonstrating the model’s strong understanding of spatial relationships. In most tasks, SpatialVLA outperforms the state-of-the-art generalist manipulation policies but struggles with long-horizon tasks in LIBERO-Long, due to the lack of architecture design for long-horizon observation.
@@
Fig. 5 summarizes the results of the Franka robot fine-tuning evaluation. In single-task tests, SpatialVLA and Diffusion Policy show similar accuracy (82% vs 81%), outperforming OpenVLA and Octo. However, in the instruction following tasks, SpatialVLA improves by +12% over OpenVLA, while Diffusion Policy struggles with a 26% success rate. In multi-tasks, SpatialVLA leverages its pre-training on OXE and 3D perception capabilities to achieve a 57% accuracy rate, surpassing other generalist policies. In summary, SpatialVLA demonstrates its versatility as a generalist robot control policy, achieving better performance across various tasks, and can be effectively used as an initialization for new robot tuning.
@@ IV-C Evaluating Spatial Understanding Capability
Evaluation Setup and Comparisons. As shown in Figs. 6 and III, we evaluate the spatial understanding capabilities of SpatialVLA through on three robot setups: Franka Robot fine-tuning, WidowX Robot zero-shot, and Libero-Spatial fine-tuning. The tasks exhibit varying spatial complexities, with the Franka task involving prompt understanding (e.g., #1 place plush toy closest to robot on car), the WidowX task featuring explicit height changes (e.g., #2 put green cup on the pink cloth), and the LIBERO-Spatial task involving object layout variations. Seven mainstream policies, namely Diffusion Policy, Octo, RT-1-X, OpenVLA, TraceVLA, and RoboVLM, are employed for comparison.
Evaluation Results. Compared to existing policies, SpatialVLA shows superior spatial understanding, achieving 73% accuracy in Franka task #1, which involves spatial prompts, and significantly improving manipulation capabilities for complex positional changes in the out-of-distribution WidowX Zero-shot tasks #2-4. Similar results are observed in the LIBERO-Spatial task suite (88.2% success rate). Policies like Octo, Diffusion Policy, and OpenVLA, which lack integrated depth information, face significant challenges in adapting to spatial layout changes, yielding a success rate consistently lower than 50%. Consequently, we suggest integrating 3D information (Sec. III), including depth or point cloud, into the VLA framework to improve the model’s adaptability and robustness in spatial layout variations.
@@
#All $[1]$. SpatialVLA 81.6% 70.7% 79.2% 85.4% 41.7% 33.3% 91.7% 87.5%
Linear Discretization $[2]$. $\sim$ linear 256 bins 40.7% 19.0% 47.1% 52.9% 41.7% 33.3% 87.5% 70.8%
Distribution $[3]$. $\sim$ uniform distribution 77.9% 28.0% 64.2% 55.0% 45.8% 12.5% 79.2% 54.2%
$[4]$. resolution 1026 74.4% 67.3% 59.1% 54.2% 45.8% 25.0% 66.7% 54.2%
Action Grids $[5]$. resolution 4610 76.7% 68.0% 69.8% 79.2% 41.7% 33.3% 83.3% 75.0%
Resolution $[6]$. resolution 6166 80.9% 74.0% 74.0% 79.2% 41.7% 33.3% 95.8% 87.5%
$[7]$. resolution 8194 81.6% 70.7% 79.2% 85.4% 41.7% 33.3% 91.7% 87.5%
Encoding $[8]$. -- ego3d encoding 68.9% 70.3% 66.7% 62.0% 54.2% 12.5% 75.0% 37.5%
@@ IV-D Ablations on Design Decisions
Pre-training in Mixture Dataset. The pre-training ablations in Tab. IV are conducted on a mixture dataset that combines Google Fractal [6] and BridgeData V2 [64]. All the models are trained from scratch on 8 A100 GPUs with 128 batch size for 120k steps. We select four tasks from the SimplerEnv benchmark [35], namely “Pick Coke Can” and “Move Near” on the Google Robot, as well as “Put Carrot on Plate” and “Put Eggplant in Yellow Basket” on the WidowX Robot, to dissect the model’s component-wise performance.
@@
Compared to 1026-resolution action grids (#1v.s.#4), where $\text{M}_{\text{trans}}=\text{M}_{\text{trans}}=512,\text{M}_{\text{grip}}=2$, SpatialVLA with 8194-resolution action grids ($\text{M}_{\text{trans}}=\text{M}_{\text{trans}}=4096,\text{M}_{\text{grip}}=2$) achieves significant performance boosts, particularly in “move near” and “put eggplant in yellow basket” tasks, with success rate increments of +31.2% and +33.3%. Additionally, we find that lower-resolution models tend to learn smaller actions, causing slow motion issues, and high-resolution models exhibit improved transfer performance in the fine-tuning stage.
According to the ablation results (#1v.s.#8), the proposed egocentric 3D position encoding (ego3d), incorporating 3D point cloud features, helps the model overcome varied lighting, color, textures, and camera poses, yielding stronger generalizability in diverse manipulation scenarios. Models w/o ego3d suffer a significant performance drop in variant aggregation, from 81.6% and 79.2% to 68.9% and 66.7%, due to their inability to adapt to scene changes. During pre-training, we also observe from (#1v.s.#9) that freezing the language embedding and sharing a trainable spatial embedding helps to improve the model’s manipulation capabilities, which is also beneficial for faster training and instruction following.
Post-training in Domain Dataset. We conduct post-training ablations in Tab. V, separately fine-tuning on large-scale datasets Google Fractal and BridgeData V2 and BridgeData V2, and comparing full fine-tuning and LoRA-tuning on the small-scale LIBERO datasets [36]. The spatial embedding adaption denotes partitioning spatial grids from the new dataset Gaussian distribution and updating the spatial feature embedding with the grids.
On large-scale datasets (#1v.s.#2), models fine-tuned with spatial embedding adaptation yield marginal gains of +2.9% in visual matching on Move Near), as the large-scale dataset distribution closely matches the pre-training distribution, allowing the model to learn fine-grained features thereby limiting the benefits of the adaption. While, on the LIBERO small dataset tasks (#4v.s.#5), initializing the feature grid with the new distribution boosts model performance by +4.6%, +5.1%, +2.2%, and +5.4% on LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long, respectively. As shown in Fig. 7, feature adaptation from the new distribution aligns pre-trained spatial features with the target fine-tuned model, improving initialization and accelerating convergence. Moreover, LoRA fine-tuning outperforms full-parameter fine-tuning on small dataset tasks (#3v.s.#4), making LoRA the preferred method for small datasets.